OpenStudy (anonymous):

solve the system of equations. enter answer as an ordered pair 3x+5y=43 5x+7y=65

OpenStudy (anonymous):

solve the first equation for x

OpenStudy (anonymous):

plug that into the 2nd equation

OpenStudy (anonymous):

then solve for y

OpenStudy (anonymous):

then plug both answers into either equation

OpenStudy (zepp):

you can solve that question by using the reduction method.

OpenStudy (anonymous):

i want the answer not how to solve it :[

OpenStudy (anonymous):

that is agains't the code of conduct

OpenStudy (anonymous):

i'm just playing, I have read the code of conduct

OpenStudy (zepp):

Here, 3x+5y=43 5x+7y=65 Let's transform that 1.4(3x+5y=43) 4.2x+7y=60.2

OpenStudy (zepp):

Now use the reduction method.